# Effortless, Simulation-Efficient Bayesian Inference using Tabular Foundation Models

### v1

@@ Abstract
Simulation-based inference (SBI) offers a flexible and general approach to performing Bayesian inference: In SBI, a neural network is trained on synthetic data simulated from a model and used to rapidly infer posterior distributions for observed data. A key goal for SBI is to achieve accurate inference with as few simulations as possible, especially for expensive simulators. In this work, we address this challenge by repurposing recent probabilistic foundation models for tabular data: We show how tabular foundation models—specifically TabPFN—can be used as pre-trained autoregressive conditional density estimators for SBI. We propose Neural Posterior Estimation with Prior-data Fitted Networks (NPE-PF) and show that it is competitive with current SBI approaches in terms of accuracy for both benchmark tasks and two complex scientific inverse problems. Crucially, it often substantially outperforms them in terms of simulation efficiency, sometimes requiring orders of magnitude fewer simulations. NPE-PF eliminates the need for inference network selection, training, and hyperparameter tuning. We also show that it exhibits superior robustness to model misspecification and can be scaled to simulation budgets that exceed the context size limit of TabPFN. NPE-PF provides a new direction for SBI, where training-free, general-purpose inference models offer efficient, easy-to-use, and flexible solutions for a wide range of stochastic inverse problems.
@@ 1 Introduction
The field of simulation-based inference (SBI), or likelihood-free inference, aims to address this challenge by enabling Bayesian inference without requiring access to likelihoods. The basic idea of SBI is to train a neural network on synthetic data simulated from the model such that it can approximate the posterior distributions. In Neural Posterior Estimation (NPE, [4, 5]), a conditional density neural network—often, a normalizing flow [6, 7] or a diffusion model [8, 9, 10]—learns the conditional distribution of parameters given simulated data $p({\bm{\theta}}\mid{\bm{x}})$. When evaluated on empirical observations ${\bm{x}}_{o}$, the network directly returns the posterior distribution $p({\bm{\theta}}\mid{\bm{x}}_{o})$. Similarly, other SBI methods use neural networks to represent likelihoods [11, 12, 13], likelihood ratios [14, 15, 16, 17], or target several properties at once [18, 19, 20]. Compared to classical Approximate Bayesian Computation (ABC, [21, 22]) or synthetic likelihoods methods [23], neural-network-based SBI methods scale better to complex, high-dimensional simulators [24]. In addition, SBI methods often amortize the computational burden: Once trained, the network provides fast inference across multiple observations.
@@
Figure 1: Illustration of NPE and NPE-PF: Both approaches use simulations sampled from the prior and simulator. In (standard) NPE, a neural density estimator is trained to obtain the posterior. In NPE-PF, the posterior is evaluated by autoregressively passing the simulation dataset and observations to TabPFN.
These limitations raise a key question: Can we eliminate the high cost associated with training and tuning, and alleviate the need to have a large number of simulations by leveraging recent advances in foundation models? Here, we answer this question by repurposing a tabular foundation model—specifically TabPFN v2 [35]—as an inference engine for SBI. TabPFN is trained to perform in-context learning on tabular data, returning a prediction when being prompted with a training dataset and a test point. For both regression and classification, TabPFN was shown to perform exceptionally well, in particular in low data regimes. Crucially, TabPFN can also serve as a density estimator by applying it in an autoregressive manner, though prior work has only explored this in limited settings [35].
Here, we show that using TabPFN autoregressively allows for the estimation of complex, high-dimensional conditional densities, making it suitable for SBI. Based on this observation, we introduce Neural Posterior Estimation with Prior-data Fitted Networks (NPE-PF), a variant of NPE that uses TabPFN as a pre-trained conditional density estimator, and therefore enables inference without any additional neural network training (Fig. 1). We show that, on benchmark tasks, NPE-PF performs competitively with other SBI approaches and often outperforms them substantially, especially for small simulation budgets. Furthermore, we leverage data-filtering schemes that enable NPE-PF to make use of large simulation budgets exceeding the context size of TabPFN and extend the approach to sequential inference settings. Finally, we show that NPE-PF remains empirically robust under model misspecification and achieves strong results on two challenging scientific applications, requiring orders of magnitude fewer simulations than standard SBI approaches.
@@ 2.1 Background
Tabular foundation models. Recent foundation models for tabular data have shown strong performance on regression and classification tasks via in-context learning. Within this work, we focus on TabPFN v2 [40, 35], a state-of-the-art transformer-based foundation model [41] for probabilistic classification and regression on tabular data. It has been pre-trained on synthetic datasets $\{(y_{i},{\bm{x}}_{i})\}_{i=0}^{M}$ with up to $M=2048$ samples from randomly generated structural causal models by maximizing the likelihood $q_{\psi}(y_{0}\mid{\bm{x}}_{0},\mathcal{T})$, where $\mathcal{T}=\{(y_{i},{\bm{x}}_{i})\}_{i=1}^{M}$ is the in-context dataset. After training, TabPFN has been tested to perform well up to a feature dimension of $d_{x}\leq 500$ and a context size of $N\leq 10^{4}$ [35].
@@ 2.2 TabPFN for simulation-based inference
where ${\mathcal{D}}^{<j}=\{\theta_{i}^{j},[{\bm{\theta}}_{i}^{<j},{\bm{x}}_{i}]\}$ with $<$$j$ denoting the vector from index $1$ up to, but not including $j$. Thus, one evaluation of the posterior $p({\bm{\theta}}\mid{\bm{x}}_{o})$ requires $d_{\theta}$ evaluations of TabPFN, making the prediction more expensive as the dimensionality increases. However, unlike standard NPE methods, this approach bypasses model fitting entirely and directly evaluates the posterior. We refer to our proposed approach as NPE-PF, in contrast to standard NPE, which requires training an inference network, e.g., a normalizing flow (Fig. 1).
@@ 2.3 Increasing the effective context size
Standard NPE can approximate arbitrarily complex conditional densities, given enough simulations and a sufficiently expressive inference network [4]. In contrast, NPE-PF is limited by TabPFN’s context size of approximately $10^{4}$ samples, beyond which additional data yields diminishing performance gains. While we expect NPE-PF to be particularly powerful for small simulation counts, it is nevertheless desirable to also be able to work with larger simulation budgets using in-context learning. Previous work has primarily tackled this problem for classification and regression settings [42]. Here, we address the problem under a broader Bayesian perspective, aiming to estimate the full posterior distribution.
@@
Filtering simulations based on relevance. To overcome the limited context size, we exploit a property of neural posterior estimation: The posterior can be estimated using only samples that are very close to the observation ${\bm{x}}_{o}$. More specifically, for any non-negative function $f_{{\bm{x}}_{o}}$ that satisfies $f_{{\bm{x}}_{o}}({\bm{x}}_{o})>0$, we can reweigh the joint distribution $p({\bm{x}},{\bm{\theta}})f_{{\bm{x}}_{o}}({\bm{x}})$ without affecting the posterior distribution at convergence (Details in Appendix Sec. B.1). This property allows us to filter simulations based on relevance without biasing the estimation. Choosing the filter as $f_{{\bm{x}}_{o}}({\bm{x}})=\mathbb{I}(d({\bm{x}},{\bm{x}}_{o})<\epsilon)$, where $d(\cdot,\cdot)$ is a suitable distance metric and $\epsilon>0$ is a threshold, effectively recovers an ABC-like selection scheme, where only simulations close to ${\bm{x}}_{o}$ are retained. In practice, $\epsilon$ is set adaptively to include the $N$ closest simulations, ensuring that the full context is used. We here consider the Euclidean distance $d({\bm{x}}_{o},{\bm{x}})=\|{\bm{x}}-{\bm{x}}_{o}\|_{2}$ after standardizing each feature dimension. This dynamic construction of the context has been previously explored and is also referred to as localization or retrieval [42, 43, 44]. This principle also underlies various SBI methods, such as the aforementioned ABC [21] or calibration kernels in NPE training [5].
@@
Filtering for simulations that are relevant or close to a given observation makes the (context) dataset dependent on that observation. For classical SBI methods, this step would require re-training the density estimator for each new observation, thus breaking amortization. However, NPE-PF remains amortized: since it is training-free, inference for new observations requires only a computationally negligible filtering step over a fixed set of simulations. Finally, filtering only works for conditional density estimation problems such as (simulation-based) Bayesian inference, which is our focus here. However, we also show how (Appendix Sec. B.3) this approach can be extended to unconditional density estimation on a larger number of samples. In the unconditional case, we partition the data space into regions that can be solved with local contexts.
@@
Embedding data. Similarly to context size, the feature size can be a limitation, particularly for high-dimensional observations. This limitation can be handled either by performing inference using human-crafted summary statistics (as commonly done in ABC) or using (pre-trained) embedding networks [5, 45]. In principle, NPE-PF allows for training an embedding network end-to-end (e.g., [46] in the context of classification), but we here investigate pretraining an embedding network independently of the density estimator (Appendix Sec. B.2).
2.4 Truncated sequential NPE-PF
@@
To reduce the number of simulations, sequential SBI methods have been developed that acquire simulations adaptively in rounds. We focus on Truncated Sequential NPE (TSNPE) [34], which mitigates bias by exploiting the (approximate) invariance of the posterior under modifications outside of its high-density regions (HDR). Formally, let $\text{HDR}^{1-\alpha}_{p}:=\{{\bm{\theta}}\mid p({\bm{\theta}}\mid{\bm{x}}_{o})>k_{\alpha}\}$, where $k_{\alpha}$ denotes the density threshold containing $\alpha\%$ of the probability mass. Then, given a truncated prior $\tilde{p}({\bm{\theta}})=\mathbb{I}({\bm{\theta}}\in\text{HDR}_{p}^{1-\alpha})p({\bm{\theta}})$, the posterior will be approximately preserved for ${\bm{x}}_{o}$. In TSNPE, a density estimator is trained in rounds. Starting from the prior, the approximation $q_{\phi}$ after the first round can then be used to approximate $\text{HDR}_{p}^{1-\alpha}\approx\text{HDR}_{q_{\phi}}^{1-\alpha}$. In subsequent rounds, the truncated prior is used to generate new simulations adaptively and refine $q_{\phi}$ through continued training. Notably, the latter becomes redundant by replacing the density estimator with the TabPFN regressor, to which we will refer as TSNPE-PF. This shifts the main computational burden to truncating the prior distribution, which can be achieved either by approximate approaches such as sampling-importance resampling or exactly by rejection sampling from the prior, i.e., checking whether a prior sample falls within the HDR. However, the latter requires repeated density estimation with the TabPFN regressor, which can be costly to do autoregressively, a problem we address next:
@@
For rejection sampling, we need to evaluate $q_{\phi}({\bm{\theta}}\mid{\bm{x}}_{o})$ for a large number of prior samples ${\bm{\theta}}\sim p({\bm{\theta}})$. In practice, the posterior is often much narrower than the prior and a large percentage of these samples will be rejected. This necessitates a fast, approximate approach for density evaluation, which we will construct using the “ratio-trick” [47]. We first sample an initial set of posterior samples $\{{\bm{\theta}}_{i}\}_{i}$ from the NPE-PF posterior once and assign these samples the class label $y=1$. We then contrast these samples with samples from a uniform distribution $\mathcal{U}({\bm{\theta}};{\bm{\theta}}_{min},{\bm{\theta}}_{max})$ to which we assign the class label $y=0$. The Bayes optimal classifier for this classification problem is given by
$P_{{\bm{x}}_{o}}(y=1\mid\theta)=\frac{q_{\phi}({\bm{\theta}}\mid{\bm{x}}_{o})}{q_{\phi}({\bm{\theta}}\mid{\bm{x}}_{o})+\mathcal{U}({\bm{\theta}};{\bm{\theta}}_{min},{\bm{\theta}}_{max})}\quad\Rightarrow\quad q_{\phi}({\bm{\theta}}\mid{\bm{x}}_{o})=\frac{P_{{\bm{x}}_{o}}(y=1\mid{\bm{\theta}})}{1-P_{{\bm{x}}_{o}}(y=1\mid{\bm{\theta}})},$
@@
within the bounds ${\bm{\theta}}_{min}$, ${\bm{\theta}}_{max}$. After autoregressively sampling from the desired posterior to construct the training dataset, this approach, which we will refer to as ratio-based density evaluation, requires only a single forward pass using the TabPFN classifier and is, therefore, significantly faster than autoregressive density evaluation, especially as the parameter dimension $d_{\theta}$ or the number of required density evaluations grows.

### v2

@@ Abstract
Simulation-based inference (SBI) offers a flexible and general approach to performing Bayesian inference: In SBI, a neural network is trained on synthetic data simulated from a model and used to rapidly infer posterior distributions for observed data. A key goal for SBI is to achieve accurate inference with as few simulations as possible, especially for expensive simulators. In this work, we address this challenge by repurposing recent probabilistic foundation models for tabular data: We show how tabular foundation models—specifically TabPFN—can be used as pre-trained autoregressive conditional density estimators for SBI. We propose Neural Posterior Estimation with Prior-data Fitted Networks (NPE-PFN) and show that it is competitive with current SBI approaches in terms of accuracy for both benchmark tasks and two complex scientific inverse problems. Crucially, it often substantially outperforms them in terms of simulation efficiency, sometimes requiring orders of magnitude fewer simulations. NPE-PFN eliminates the need for selecting and training an inference network and tuning its hyperparameters. We also show that it exhibits superior robustness to model misspecification and can be scaled to simulation budgets that exceed the context size limit of TabPFN. NPE-PFN provides a new direction for SBI, where training-free, general-purpose inference models offer efficient, easy-to-use, and flexible solutions for a wide range of stochastic inverse problems.
@@ 1 Introduction
The field of simulation-based inference (SBI), or likelihood-free inference, aims to address this challenge by enabling Bayesian inference without requiring access to likelihoods. The basic idea of SBI is to train a neural network on synthetic data simulated from the model such that it can approximate the posterior distributions. In Neural Posterior Estimation (NPE, [4, 5]), a conditional density neural network—often, a normalizing flow [6, 7] or a diffusion model [8, 9, 10]—learns the conditional distribution of parameters given simulated data $p({\bm{\theta}}\mid{\bm{x}})$. When evaluated on empirical observations ${\bm{x}}_{o}$, the network directly returns the posterior distribution $p({\bm{\theta}}\mid{\bm{x}}_{o})$. Similarly, other SBI methods use neural networks to represent likelihoods [11, 12, 13], likelihood ratios [14, 15, 16, 17], or target several properties at once [18, 19, 20, 21]. SBI has been used for scientific discovery in various domains, including astrophysics [22] and neuroscience [23].
Compared to classical Approximate Bayesian Computation (ABC, [24, 25]) or synthetic likelihoods methods [26], neural-network-based SBI methods scale better to complex, high-dimensional simulators [27]. In addition, SBI methods often amortize the computational burden: Once trained, the network provides fast inference across multiple observations.
@@
Figure 1: Comparison of NPE to NPE-PFN (ours): Both approaches use simulations sampled from the prior and simulator. In (standard) NPE, a neural density estimator is trained to obtain the posterior. In NPE-PFN, the posterior is evaluated by autoregressively passing the simulation dataset and observations to TabPFN.
These limitations raise a key question: Can we eliminate the high cost associated with training and tuning, and alleviate the need to have a large number of simulations by leveraging recent advances in foundation models? Here, we answer this question by repurposing a tabular foundation model—specifically Tabular Prior-data Fitted Networks v2 (TabPFN) [38]—as an inference engine for SBI. TabPFN is trained to perform in-context learning on tabular data (i.e., structured data organized in rows and columns), returning a prediction when being prompted with a training dataset and a test point. For both regression and classification, TabPFN was shown to perform exceptionally well, in particular in low data regimes. Crucially, TabPFN can also serve as a density estimator by applying it in an autoregressive manner, though prior work has only explored this in limited settings [38].
Here, we show that using TabPFN autoregressively allows for the estimation of complex, high-dimensional conditional densities, making it suitable for SBI. Based on this observation, we introduce Neural Posterior Estimation with Prior-data Fitted Networks (NPE-PFN), a variant of NPE that uses TabPFN as a pre-trained conditional density estimator, and therefore enables inference without any additional neural network training (Fig. 1). On benchmark tasks, NPE-PFN performs competitively with other SBI approaches and often outperforms them substantially, especially for small simulation budgets. Furthermore, we leverage data-filtering schemes that enable NPE-PFN to make use of large simulation budgets exceeding the context size of TabPFN and extend the approach to sequential inference settings. Finally, we show that NPE-PFN remains empirically robust under model misspecification and achieves strong results on two challenging scientific applications, requiring orders of magnitude fewer simulations than standard SBI approaches.
@@ 2.1 Background
Tabular foundation models. Recent foundation models for tabular data have shown strong performance on regression and classification tasks via in-context learning, in which predictions for new inputs are made by conditioning the model on a context dataset [43]. We focus on Prior-data Fitted Networks (PFNs) [44, 45], specifically TabPFN [38], a state-of-the-art transformer-based PFN foundation model [46] for probabilistic classification and regression on tabular data. It has been pre-trained on synthetic datasets $\{(y_{i},{\bm{x}}_{i})\}_{i=0}^{M}$ with up to $M=2048$ samples from randomly generated structural causal models by maximizing the likelihood $q_{\psi}(y_{0}\mid{\bm{x}}_{0},\mathcal{T})$, where $\mathcal{T}=\{(y_{i},{\bm{x}}_{i})\}_{i=1}^{M}$ is the in-context dataset. After training, TabPFN has been tested to perform well up to a feature dimension of $d_{x}\leq 500$ and a context size of $N\leq 10^{4}$ [38].
@@ 2.2 TabPFN for simulation-based inference
where ${\mathcal{D}}^{<j}=\{\theta_{i}^{j},[{\bm{\theta}}_{i}^{<j},{\bm{x}}_{i}]\}$ with <<$j$ denoting the vector from index $1$ up to, but not including $j$. Thus, using an arbitrary but fixed order of the parameter dimensions, one evaluation of the posterior $p({\bm{\theta}}\mid{\bm{x}}_{o})$ requires $d_{\theta}$ evaluations of TabPFN, making the prediction more expensive as the dimensionality increases. However, unlike standard NPE methods, this approach bypasses model fitting entirely and directly evaluates the posterior. We refer to our proposed approach as NPE-PFN, in contrast to standard NPE, which requires training an inference network, e.g., a normalizing flow (Fig. 1, pseudocode for NPE-PFN in Appendix Alg. 1).
@@ 2.3 Increasing the effective context size
Standard NPE can approximate arbitrarily complex conditional densities, given enough simulations and a sufficiently expressive inference network [4]. In contrast, NPE-PFN is limited by TabPFN’s context size of approximately $10^{4}$ samples, beyond which additional data yields diminishing performance gains. While we expect NPE-PFN to be particularly powerful for small simulation counts, it is nevertheless desirable to also be able to work with larger simulation budgets using in-context learning.
@@
Filtering simulations based on relevance. To overcome the limited context size, previous work proposed to use a subset of the training data containing only the nearest neighbors of a given test point as the context dataset [47]. Here, we show that this procedure, and generalizations of it, are sound from a Bayesian perspective when estimating the full posterior distribution. To this end, we exploit a key property of (neural) posterior estimation: The posterior can be estimated using only samples that are very close to the observation ${\bm{x}}_{o}$. More specifically, for any non-negative function $f_{{\bm{x}}_{o}}$ that satisfies $f_{{\bm{x}}_{o}}({\bm{x}}_{o})>0$, we can reweigh the joint distribution $p({\bm{x}},{\bm{\theta}})f_{{\bm{x}}_{o}}({\bm{x}})$ without affecting the posterior distribution at convergence (details in Appendix Sec. B.2). This property allows us to filter simulations based on relevance without biasing the estimation. Choosing the filter as $f_{{\bm{x}}_{o}}({\bm{x}})=\mathbb{I}(d({\bm{x}},{\bm{x}}_{o})<\epsilon)$, where $d(\cdot,\cdot)$ is a suitable distance metric and $\epsilon>0$ is a threshold, effectively recovers an ABC-like selection scheme, where only simulations close to ${\bm{x}}_{o}$ are retained. In practice, $\epsilon$ is set adaptively to include the $N_{\mathrm{filter}}$ closest simulations, ensuring that the full context is used. Here, we consider the Euclidean distance $d({\bm{x}}_{o},{\bm{x}})=\|{\bm{x}}-{\bm{x}}_{o}\|_{2}$ after standardizing each feature dimension. As noted above, this filtering based on nearest neighbors has previously been explored and is also referred to as localization or retrieval [47, 48, 49]. This filtering principle also underlies various SBI methods, such as the aforementioned ABC [24] or calibration kernels in NPE training [5].
@@
Filtering for simulations that are relevant or close to a given observation makes the (context) dataset dependent on that observation. For classical SBI methods, this step would require re-training the density estimator for each new observation, thus breaking amortization. However, NPE-PFN remains amortized: Since it is training-free, inference for new observations requires only a computationally negligible filtering step over a fixed set of simulations. Finally, filtering only works for conditional density estimation problems such as (simulation-based) Bayesian inference, which is our focus here. However, we also show how (Appendix Sec. B.4) this approach can be extended to unconditional density estimation on a larger number of samples. In the unconditional case, we partition the data space into regions that can be solved with local contexts.
@@
Embedding data. Similarly to context size, the feature size can be a limitation, particularly for high-dimensional observations. This limitation can be handled either by performing inference using human-crafted summary statistics (as commonly done in ABC) or using (pre-trained) embedding networks [5, 50]. In principle, NPE-PFN allows for training an embedding network end-to-end (e.g., [51] in the context of classification), but we here investigate pre-training an embedding network independently of the density estimator (Appendix Sec. B.3).
2.4 Truncated sequential NPE-PFN
@@
To reduce the number of simulations, sequential SBI methods have been developed that acquire simulations adaptively in rounds for one observation ${\bm{x}}_{o}$. We focus on Truncated Sequential NPE (TSNPE) [37], which mitigates bias by exploiting the (approximate) invariance of the posterior under modifications outside of its high-density regions (HDR). Formally, let $\text{HDR}^{1-\alpha}_{p}:=\{{\bm{\theta}}\mid p({\bm{\theta}}\mid{\bm{x}}_{o})>k_{\alpha}\}$, where $k_{\alpha}$ denotes the density threshold containing $\alpha\%$ of the probability mass. Then, given a truncated prior $\tilde{p}({\bm{\theta}})=\mathbb{I}({\bm{\theta}}\in\text{HDR}_{p}^{1-\alpha})p({\bm{\theta}})$, the posterior will be approximately preserved for ${\bm{x}}_{o}$. In TSNPE, a density estimator is trained in rounds. Starting with the full prior in the first round, the trained posterior approximation $q({\bm{\theta}}\mid{\bm{x}}_{o})$ is used to estimate $\text{HDR}_{p}^{1-\alpha}\approx\text{HDR}_{q}^{1-\alpha}$ and to truncate the prior to the HDR of the posterior approximation for the next round. In each subsequent round, the newly truncated prior is used to generate new simulations adaptively and refine $q({\bm{\theta}}\mid{\bm{x}}_{o})$ through continued training. Notably, training becomes redundant by replacing the density estimator with the TabPFN regressor, to which we will refer as TSNPE-PFN. In this training-free setting, the main computational burden shifts to truncating the prior distribution. Prior truncation can be achieved either by approximate approaches, such as sampling-importance resampling, or exactly by rejection sampling from the prior, i.e., checking whether a prior sample falls within the HDR. However, the latter approach requires repeated density estimation with the TabPFN regressor to evaluate $q({\bm{\theta}}\mid{\bm{x}}_{o})$, which can be costly to do autoregressively, a problem we address next.
@@
For rejection sampling, we need to evaluate $q({\bm{\theta}}\mid{\bm{x}}_{o})$ for a large number of prior samples ${\bm{\theta}}\sim p({\bm{\theta}})$. In practice, the posterior is often much narrower than the prior and a large percentage of these samples will be rejected. This high rejection rate necessitates a fast, approximate approach for density evaluation, which we will construct using the “ratio-trick” [52]. We first sample an initial set of posterior samples $\{{\bm{\theta}}_{i}\}_{i}$ from the NPE-PFN posterior once and assign these samples the class label $y=1$. We then contrast these samples with samples from a uniform distribution $\mathcal{U}({\bm{\theta}};{\bm{\theta}}_{min},{\bm{\theta}}_{max})$ to which we assign the class label $y=0$. The Bayes optimal classifier for this classification problem is given by
$P_{{\bm{x}}_{o}}(y=1\mid{\bm{\theta}})=\frac{p({\bm{\theta}}\mid{\bm{x}}_{o})}{p({\bm{\theta}}\mid{\bm{x}}_{o})+\mathcal{U}({\bm{\theta}};{\bm{\theta}}_{min},{\bm{\theta}}_{max})},\quad\text{thus,}\quad p({\bm{\theta}}\mid{\bm{x}}_{o})=\frac{P_{{\bm{x}}_{o}}(y=1\mid{\bm{\theta}})}{1-P_{{\bm{x}}_{o}}(y=1\mid{\bm{\theta}})},$
@@
within the bounds ${\bm{\theta}}_{min}$, ${\bm{\theta}}_{max}$. After autoregressively sampling from the desired posterior to construct the training dataset, this approach, which we will refer to as ratio-based density evaluation, requires only a single forward pass using the TabPFN classifier and is, therefore, significantly faster than autoregressive density evaluation, especially as the parameter dimension $d_{\theta}$ or the number of required density evaluations grows (pseudocode for TSNPE-PFN in Appendix Alg. 2).